# LLoCO: Learning Long Contexts Offline

### v1

@@ 4.2 Ablation Study
Finetuning LLaMA with Original Context One interesting question is how well LoRA finetuning works over original uncompressed context, and how does that compare to LLoCO’s approach. To investigate this, we conduct additional finetuning experiments on LLaMA-7B-4k/32k, where we append the original uncompressed context as a system prompt during finetuning. These models are finetuned following the same setup as LLoCO.
@@
LLaMA-7B-4k 40.45 16.67 14.62 14.42 32.47 23.44
LLaMA-7B-32k 38.88 21.72 14.58 16.76 31.58 24.70
LLaMA-7B-4k w. Finetuning 35.00 17.80 15.49 21.41 41.89 26.32 (+2.88%)
LLaMA-7B-32k w. Finetuning 40.12 29.71 16.36 28.72 41.68 31.32 (+6.62%)
@@
Setup QuA QAS QM NQA HQA Avg.
LLoCO w. Separate Finetuning 41.51 29.01 16.68 28.34 37.82 30.67
LLoCO w. Combined Finetuning 47.07 24.95 12.77 27.93 35.55 29.65
@@ 4.3 Evaluation on LongBench
We further evaluate LLoCO on LongBench (Bai et al., 2023), which consists of 6 subtasks. Given that the primary applications of LLoCO are document question answering and summarization, our evaluation focuses on the SingleDoc QA, MultiDoc QA, and Summarization tasks. These three tasks overlap with some datasets (e.g. Qasper, QMSum) we evaluated in Section 4.1. However, the validation sets differ as LongBench samples a specific subset of examples for each dataset. We have rigorously ensured that LongBench does not feature any questions that overlap with those used in the LLoCO training, thereby completely removing any risk of data leakage. To achieve the best performance of LLoCO, we choose LoRA adaptors tailored for each particular category/dataset in our evaluation.
Specifically, for the SingleDoc tasks, we use the NQA LLoCO and Qasper LLoCo for the NQA and QAS tasks, respectively. For MultiField-En (MFE)  (Bai et al., 2023), we use GPT-4-turbo to generate question-answering pairs from the training documents and finetune LLoCO using these pairs. For MultiDoc tasks, we use the combined finetuned LLoCO from Section 2, which works well for all tasks. For the Summarization tasks, we create a training dataset by combining the entire QMSum (Zhong et al., 2021) training data with 5,000 randomly sampled examples from the MultiNews (Fabbri et al., 2019) training data. We then finetune LLoCO on this combined dataset.
@@
Our evaluation results show that LLoCO outperforms the LLaMA2 baseline on 5 out of the 9 datasets. In particular, LLoCO excels in the MultiDoc QA task, achieving significantly better results on all three datasets. LLoCO also demonstrates comparable performance in two datasets (MultiNews, Qasper), but falls short in the GovReport Huang et al. (2021) and MultiField-En datasets. The GovReport dataset is a summarization task that requires the model to generate a one-page summary. We found that LLoCO does not perform well when the generated content is lengthy, which is a limitation of our approach. The lower performance on the MultiField-En dataset could be attributed to the data being out-of-distribution compared to our training data, as this dataset does not provide any training examples. Despite these limitations, the average score of LLoCO across all datasets is higher than that of the LLaMA2 baseline, demonstrating the overall effectiveness of our approach.
@@ 4.4 Needle In A Haystack
We further investigate LLoCO’s proficiency to retrieve information across different positions of context window with varying lengths using the renowned Needle In A Haystack task (gkamradt, 2023). Tailoring this to our pipeline, we randomly select a long article exceeding 32k tokens from the NarrativeQA dataset as the “haystack”. This article is used to test the LLoCO model, which has been finetuned on this dataset as discussed in § 4.1.
@@
(a) LLaMA2-7B-32k
(b) LLoCO
Figure 3: Fixed needle retrieval task. The sampled article (”haystack”) starts with ”Mary, ..., a gentle, fashionable girl...”, and a context-relevant needle was curated as ”Mary’s favorite fashion designer was Coco Chanel when she was a teenager. Who was Mary’s favorite fashion designer when she was a teenager?
@@
4.5 Inference Latency & Finetuning Throughput
@@
Additionally, we assess the finetuning throughput on the NarrativeQA dataset, which mainly consists of samples exceeding $32$k tokens. We compare the 7B $32$k baseline finetuned on 8 A100s and our LLoCO finetuned on both 8 A100s and 8 A6000s, all with a per-device batch size of 1 and a global batch size of 32, using 4 gradient accumulation steps. Figure 5 shows that our LLoCO achieves 11.52×11.52$\times$11.52 × and 6.82×6.82$\times$6.82 × throughput on A100s and A6000s, respectively. This highlights that LLoCO not only achieves competitive results, but also improves performance with much greater efficiency compared to finetuning the baseline with full context (as shown in § 4.2).
Figure 5: Left: End-to-end decoding per-token latency (ms) on A100 and A6000 GPUs. LLaMA2 without compression runs out of VRAM for sequences of 64k and 128k on A100, as well as for 32k sequences on A6000. Right: Train samples per second when finetuning on NarrativeQA for LLaMA2-7B-32k without compression and LLoCO.
@@ 5 Conclusion
We propose LLoCO, a new paradigm that addresses the long-context challenge by preprocessing the context before inference through parameter-efficient finetuning. Our approach extends the effective context window of a LLaMA2-7B model with a 4k tokens context size to handle up to 128k tokens. Evaluations on several long-context question-answering datasets show that LLoCO significantly outperforms in-context learning while using 30 times fewer tokens during inference, making it a promising solution for efficient long context processing.
@@
Appendix A Appendix
A.1 Extended Experimental Settings
In this part, we provide details of our experimental settings in the main text. For the experiments presented in Sec 4.1, we summarize the hyperparameter settings in Table 5. We use all samples from the train split of each dataset for finetuning, except for HotpotQA, from which we randomly select $20,000$ samples from its train split.
@@ A.1 Extended Experimental Settings
A.2 More Details on Datasets

### v2

@@ 4.2 Ablation Study
In this section, we present ablation studies comparing LLoCO under various setups, including finetuning LLaMA with the original context and evaluating different compression ratios. Additional studies, such as exploring alternative context encoders, are provided in Appendix D.
Finetuning LLaMA with Original Context One interesting question is how well LoRA finetuning works over original uncompressed context, and how does that compare to LLoCO’s approach. To investigate this, we conduct additional finetuning experiments on LLaMA2-7B-4k/32k, where we append the original uncompressed context as a system prompt during finetuning. These models are finetuned following the same setup as LLoCO. We do not explore the impact of finetuning LLaMA2-7B-128k with uncompressed context due to constraints on compute resources.
@@
Exploring the Impact of Compression Ratios In this ablation study, we evaluate LLoCO under various compression ratios to analyze their effect on performance. The original LLoCO setup, adapted from AutoCompressor, applies a 30x compression ratio, reducing 1536 tokens to 50 tokens. To explore the impact of different ratios, we also compress 1024 tokens and 2048 tokens into 50 tokens, corresponding to compression ratios of 20x and 40x, respectively. We assess these configurations on the QuALITY, QMSum, and NarrativeQA datasets, with the results presented in Figure 2.
Figure 2: Impact of compression ratio on LLoCO’s performance.
Our findings suggest that performance remains relatively stable across different compression ratios, with both 20x and 30x ratios consistently outperforming 40x across all datasets. Interestingly, the 30x ratio performs on par with, and occasionally slightly better than, the 20x ratio. We attribute this to the context encoder (AutoCompressor) being originally optimized for the 30x setting, allowing it to maintain high performance even at higher compression rates.
@@
LLaMA2-7B-4k 40.45 16.67 14.62 14.42 32.47 23.44
LLaMA2-7B-32k 38.88 21.72 14.58 16.76 31.58 24.70
LLaMA2-7B-4k w. Finetuning 35.00 17.80 15.49 21.41 41.89 26.32 (+2.88%)
LLaMA2-7B-32k w. Finetuning 40.12 29.71 16.36 28.72 41.68 31.32 (+6.62%)
@@ 4.3 Evaluation on LongBench
We further evaluate LLoCO on LongBench (Bai et al., 2023), which consists of 6 subtasks. Given that the primary applications of LLoCO are document question answering and summarization, our evaluation focuses on the SingleDoc QA, MultiDoc QA, and Summarization tasks. There exists overlap with some datasets (e.g. Qasper, QMSum) we evaluated in Section 4.1. However, the validation sets differ as LongBench samples a specific subset of examples for each dataset. We have rigorously ensured that LongBench does not feature any questions that overlap with those used in the LLoCO training, thereby completely removing any risk of data leakage. To achieve the best performance of LLoCO, we choose LoRA adaptors tailored for each particular category/dataset in our evaluation.
Specifically, for the SingleDoc tasks, we use the NQA LLoCO and Qasper LLoCo for the NQA and QAS tasks, respectively. For MultiField-En (MFE)  (Bai et al., 2023), we use GPT-4-turbo to generate question-answering pairs from the training documents and finetune LLoCO using these pairs. For MultiDoc tasks, we use the combined finetuned LLoCO from Section D, which works well for all tasks. For the Summarization tasks, we create a training dataset by combining the entire QMSum (Zhong et al., 2021) training data with 5,000 randomly sampled examples from the MultiNews (Fabbri et al., 2019) training data. We then finetune LLoCO on this combined dataset.
@@
Our evaluation results show that LLoCO outperforms the baselines on 5 out of the 9 datasets. In particular, LLoCO excels in the MultiDoc QA task, achieving much better results on all three datasets. LLoCO also demonstrates comparable performance in two datasets (MultiNews, Qasper), but falls short in the GovReport Huang et al. (2021) and MultiField-En datasets. The GovReport dataset is a summarization task that requires the model to generate a one-page summary. We found that LLoCO does not perform well when the generated content is lengthy, which currently is a limitation of our approach. The lower performance on the MultiField-En dataset could be attributed to the data being out-of-distribution compared to our training data, as this dataset does not provide any training examples. Despite these limitations, the average score of LLoCO across all datasets is higher than that of the LLaMA2 baseline, highlighting the overall effectiveness of our approach.
@@ 4.4 Needle In A Haystack
We further investigate LLoCO’s proficiency to retrieve information across different positions of context window with varying lengths using the renowned Needle In A Haystack task (gkamradt, 2023). Tailoring this to our pipeline, we randomly select a long article exceeding 32k tokens from the NarrativeQA dataset as the “haystack". This article is used to test the LLoCO model, which has been finetuned on this dataset as discussed in Section 4.1.
@@
Figure 3: Fixed needle retrieval task. The sampled article ("haystack") starts with "Mary, ..., a gentle, fashionable girl...", and a context-relevant needle was curated as "Mary’s favorite fashion designer was Coco Chanel when she was a teenager. Who was Mary’s favorite fashion designer when she was a teenager?"
@@
4.5 Inference Latency
@@
We additionally measure LLoCO’s finetuning throughput improvement over finetuning LLaMA with the original context, where LLoCO achieves a up to 11.52×11.52$\times$11.52 × higher throughput. The detailed results can be found in Appendix C.
Figure 5: End-to-end decoding per-token latency (ms) on A100 and A6000 GPUs. LLaMA2 without compression runs out of VRAM for sequences of 64k and 128k on A100, and for 32k sequences on A6000.
@@ 5 Conclusion
We propose LLoCO, a new paradigm that addresses the long-context challenge by preprocessing the context before inference through parameter-efficient finetuning. Our approach extends the effective context window of a LLaMA2-7B model with a 4k tokens context size to handle up to 128k tokens. Evaluations on extensive long-context benchmarks show that LLoCO significantly outperforms in-context learning while using 30×30$\times$30 × fewer tokens during inference, making it a promising solution for efficient long-context processing.
6 Limitations
One limitation of our work is that the context encoder is tied to a specific LLM model. It compresses the original context into a representation aligned with the model. Therefore, for a different model, a separate context encoder must be trained to ensure alignment. Training a context encoder such as AutoCompressor for LLaMA2-7B is costly, requiring about 15 billion tokens of data. On the other hand, studies such as Morris et al. (2023) suggest that original texts can be recovered from sentence embeddings, indicating the possibility of constructing context embeddings that are easily and universally recoverable. Developing a model-agnostic context encoder that can adapt to various models without additional training poses an important challenge and a promising area for future research.
In LLoCO’s pipeline, we separate documents into groups and finetune a LoRA adapter for each group. In a standard RAG pipeline, top $k$ documents relevant to the user’s query are retrieved. However, these $k$ documents may come from different groups, each corresponding to a different LoRA adapter, and we can apply only one adapter at a time, which is a limitation. An intriguing future direction would be to investigate the composition of multiple LoRA adapters simultaneously to further enhance QA performance.
There are differences in the effectiveness of LLoCO across various datasets and tasks. As shown in Section 4.3 and 4.4, our performance on GovReport and MultiField-En does not surpass the baselines, and we do not achieve perfect results on the Needle-In-A-Haystack task. We attribute these issues to the limitations of the training data quality and the capabilities of the context encoder and LLaMA2 base model. We believe that improving these components will address these shortcomings.
@@ 6 Limitations
Appendix A Extended Experimental Settings
In this part, we provide details of our experimental settings in the main text. For the experiments presented in Sec 4.1, we summarize the hyperparameter settings in Table 5. We use all samples from the train split of each dataset for finetuning, except for HotpotQA, from which we randomly select $20,000$ samples from its train split. All finetuning experiments are conducted with either 4 or 8 NVIDIA RTX A6000 GPUs or A100-80G-SXM GPUs.
@@ Appendix A Extended Experimental Settings
Appendix B More Details on Datasets
@@
We additionally provide the statistics of these five datasets in Table 6.
QuA QAS QM NQA HQA
# of samples 2086 1726 272 5878 7405
# of unique articles 115 273 35 115 7306
Avg word count 4122 3572 10422 51925 952
Max word count 5967 14640 24573 353480 2694
Max token length 9953 24205 34478 556249 4000
Table 6: Statistics of the datasets. The maximum token length is counted using LLaMA2-7B’s tokenizer.
Appendix C Finetuning Throughput
We assess the finetuning throughput for both LLoCO and LLaMA2-7B-32k with the original context on the NarrativeQA dataset, which mainly consists of samples exceeding $32$k tokens. We compare the 7B $32$k baseline finetuned on 8 A100s and our LLoCO finetuned on both 8 A100s and 8 A6000s, all with a per-device batch size of 1 and a global batch size of 32, using 4 gradient accumulation steps. For the 7B $32$k baseline finetuning, samples of more than $32$k tokens were truncated to $32$k. Figure 5 shows that our LLoCO achieves 11.52×11.52$\times$11.52 × and 6.82×6.82$\times$6.82 × throughput on A100s and A6000s, respectively. This highlights that LLoCO not only achieves competitive results, but also improves performance with much greater efficiency compared to finetuning the baseline with full context (as shown in Section 4.2).
Figure 6: Samples per second when finetuning on NarrativeQA for LLaMA2-7B-32k without compression and LLoCO.
Appendix D Additional Ablation Study
Setup QuA QAS QM NQA HQA Avg.
LLoCO w. Separate Finetuning 41.51 29.01 16.68 28.34 37.82 30.67
LLoCO w. Combined Finetuning 47.07 24.95 12.77 27.93 35.55 29.65
Exploring other Context Encoders
Setup QuA QAS QM NQA
LLaMA2-7B-4k 40.45 16.67 14.62 14.42
ICAE 36.5 16.14 15.69 17.38
LLoCO 41.51 29.01 16.68 28.34
Table 8: Comparison of LLoCO with ICAE Ge et al. (2023)
By default, LLoCO utilizes AutoCompressor as its context encoder. However, LLoCO offers flexibility in its choice of context encoders and is not limited to a single method. In this ablation study, we evaluate LLoCO using another state-of-the-art context compression method, ICAE Ge et al. (2023), as the context encoder for LLoCO. Since ICAE was trained on contexts up to 4096 tokens, we truncated the inputs to this length, resulting in compressed embeddings of 128 tokens, following the approach outlined in the original paper.
Our results in Table 8 show that ICAE performs comparably to LLaMA2-7B-4K across benchmarks, while achieving a 40x compression ratio. This highlights that our proposed pipeline, which combines context compression with PEFT, is not restricted to AutoCompressor and can be generalized to other context encoders.
However, there remains a notable performance gap between ICAE and LLoCO. A key limitation of ICAE is its inability to extend the context window beyond 4K tokens, making it a less suitable option compared to AutoCompressor as the context encoder for our approach.
Appendix E LLoCO’s Serving Stage
Figure 7: The serving stage of LLoCO’s pipeline.
In Figure 7, we illustrate LLoCO’s serving pipeline. In a standard RAG, the retriever computes a text embedding for each passage, and use that as the key to retrieve passages relevant to the user query. We use the same text embedding as key, but instead retrieve the compressed token embeddings of the passages, and prepend it to the decoder LLM. Additionally, the system searches for the corresponding LoRA adaptor in the database and applies it to the decoder LLM. Applying a LoRA adaptor incurs minimal overhead to the system cost  (Hu et al., 2022). By leveraging recent work on serving multiple LoRA adaptors (Sheng et al., 2023a; Chen et al., 2023a), we can parallelize LoRA adaptor serving and simultaneously serve requests for documents associated with thousands of LoRA adaptors on a single GPU.
In our current system, we assume that all passages retrieved from the vector database for a given user query belong to the same document group, allowing us to use a single dedicated LoRA adaptor. However, LLoCO can be extended to handle cases where the retrieved passages span multiple document groups. For instance, it is feasible to design algorithms to dynamically select the most relevant LoRA adaptor based on the majority of the retrieved documents or weight the contributions of different adaptors by their relevance to the query (Muqeeth et al., 2024). Another interesting and orthogonal research direction is to explore composing multiple LoRA modules together (Huang et al., 2023), and we leave it as future work.
Appendix F Comparison with Concurrent Work
CEPE Yen et al. (2024) and SnapKV Li et al. (2024) are two concurrent work that also proposes methods for context compression. We provide a comparison with them in this section.
Setup QuA QAS QM (ROUGE-L) NQA
CEPE (official) 30.2 20.5 19.6 21.9
CEPE (reproduced) 26.40 20.54 19.38 20.33
LLoCO 41.51 29.01 20.92 28.34
Table 9: Comparison of LLoCO with CEPE Yen et al. (2024)
SingleDoc MultiDoc Summarization
NQA QAS MFE HQA WMQA MSQ Gov QMS MNews Avg.
SnapKV 20.7 29.3 42.2 34.0 24.9 14.2 28.6 23.1 26.5 27.1
LLoCO 23.1 26.1 26.3 46.2 35.6 27.3 17.6 23.4 25.0 27.8
Table 10: Comparison between LLoCO and SnapKV Li et al. (2024) on LongBench (THUDM, 2023)
Comparison with CEPE
Similar to LLoCO, CEPE requires training an encoder to generate embeddings for the context. During inference, CEPE uses cross-attention to integrate these embeddings into the original LLM decoder. The main difference is that CEPE applies the encoder at runtime, allowing it to process context in parallel, thus speeding up the inference process. In contrast, LLoCO performs context encoding offline, retrieving the preprocessed embeddings during inference.
To compare with CEPE, we evaluate their model (based on LLaMA2-7B-Chat) on QuaLITY, Qasper, QMSum, and NarrativeQA. We present two sets of results: the official evaluation performance provided by CEPE, and our reproduced results, obtained by running their publicly available code on these datasets.
As shown in Table 9, LLoCO outperforms CEPE across all datasets. Note that for QMSum dataset, CEPE provided the ROUGE-L score, while LLoCO provides the geometric mean of ROUGE-1, ROUGE-2, and ROUGE-L scores as our metric. We take the ROUGE-L score of LLoCO here for a fair comparison.
Comparison with SnapKV
Compared to LLoCO and CEPE Yen et al. (2024), which requires first finetune a encoder beforehand, SnapKV is a KV cache compression approach that is finetuning free. The key insight behind SnapKV is that only a small subset of tokens in the prompt is crucial (those that other tokens attend to), allowing the less important tokens to be discarded.
Since both LLoCO and SnapKV provide evaluation results on LongBench, we offer a side-by-side comparison of their performance. We extract SnapKV’s results using LLaMa2-7B-32k (LongChat) as the base model and compare them with ours. SnapKV provides three settings (1024, 2048, and 4096 tokens), and we use their best results for comparison.
Table 10 illustrates that overall, LLoCO and SnapKV show comparable performance across tasks. LLoCO outperforms SnapKV on five datasets, while SnapKV leads on four. Unlike SnapKV, which compresses input tokens to fixed lengths (1024, 2048, 4096), LLoCO applies a consistent 30x compression ratio, enabling the compression of up to 128k tokens into 4096 tokens. This generally results in a higher compression ratio than SnapKV on LongBench.